OpenStudy (anonymous):

convert fraction to percent: 40/50, 3/20, 12/20, 21/25, 23/25, 13/20 and 19/25 answer all the question which u know the answer thankyou

OpenStudy (anonymous):

Divide the numerator by the denominator. You'll end up with a decimal value: 40/50 = 0.8 Then change it into percent. 0.8 = 80%

OpenStudy (anonymous):

Please do not give out straight answers. It is against the Code of Conduct.

OpenStudy (anonymous):

or one more method is see the denominator and see with what no u have to multiply it to make a 100...multiply numarator with the same no....

OpenStudy (anonymous):

sure...

OpenStudy (anonymous):

That way is much easier^. Just remember that you will not always have denominators that evenly go into 100.

OpenStudy (aravindg):

@raje_raviraj I also repeat @jazy 's words" Please do not give out straight answers. It is against the Code of Conduct." http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

here it is and it is faster way to get ans...don't u think...?

OpenStudy (aravindg):

and @raje_raviraj and @Nana44255 \[\huge \color{green}{\text{WELCOME TO OPENSTUDY!}}\]

OpenStudy (anonymous):

but i agree with you...

OpenStudy (anonymous):

thanks arvind

OpenStudy (anonymous):

wow thx to everyone

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

and arvindG u like pokemon? and im from japan where pokemon was made

OpenStudy (aravindg):

wow great pikachu say hi to nana "Pika ...Pika..."

OpenStudy (aravindg):

take time to update your profile by clicking settings on top right corner both of you :)

OpenStudy (anonymous):

and can someone tell me the answer of 19/25? plz i hv many other works to do

OpenStudy (anonymous):

\[\frac{ 19 }{ 25 } \times 4 = \frac{ ? }{ 100 }\]

OpenStudy (anonymous):

Multiply 19 by 4. That is your percentage.

OpenStudy (anonymous):

OHH thankyou!!!

OpenStudy (anonymous):

You're welcome. (: